Solve the system shown. Put your answer in (x, y) format with no spaces.
2x + 9y = 0
3x + 5y = 17

Answer by oberobic(2304)

2x + 9y = 0
3x + 5y = 17
.
multiply the first equation by 3 and the second by 2; then subtract
.
6x + 27y = 0
6x + 10y = 34
--------------
17y = -34
.
y = -2
.
substitute to find 'x'
.
2x + 9(-2) = 0
2x -18 = 0
2x = 18
x = 9
.
check by substituting these values into the other equation
.
3(9) + 5(-2) = 17 ??
27 -10 = 17
correct
.
Answer: (9,-2)
.
Done.